# MaeFuse: Transferring Omni Features with Pretrained Masked Autoencoders for Infrared and Visible Image Fusion via Guided Training

### v1

@@ Abstract
In this research, we introduce MaeFuse, a novel autoencoder model designed for infrared and visible image fusion (IVIF). The existing approaches for image fusion often rely on training combined with downstream tasks to obtain high-level visual information, which is effective in emphasizing target objects and delivering impressive results in visual quality and task-specific applications. MaeFuse, however, deviates from the norm. Instead of being driven by downstream tasks, our model utilizes a pretrained encoder from Masked Autoencoders (MAE), which facilities the omni features extraction for low-level reconstruction and high-level vision tasks, to obtain perception friendly features with a low cost. In order to eliminate the domain gap of different modal features and the block effect caused by the MAE encoder, we further develop a guided training strategy. This strategy is meticulously crafted to ensure that the fusion layer seamlessly adjusts to the feature space of the encoder, gradually enhancing the fusion effect. It facilitates the comprehensive integration of feature vectors from both infrared and visible modalities, preserving the rich details inherent in each. MaeFuse not only introduces a novel perspective in the realm of fusion techniques but also stands out with impressive performance across various public datasets.
1 Introduction
@@
Multimodal sensing technology has significantly contributed to the widespread use of multimodal imaging in various fields, and the fusion of infrared and visible modalities is the most common application technique. Infrared and visible image fusion (IVIF) can merge the complementary information of these two modalities. Infrared images display thermal contours and are unaffected by lighting effects, but lack texture and color accuracy. Visible images can capture texture but depend on lighting. Fused images can combine their strengths, and thus enhancing visual clarity and the effectiveness of downstream visual tasks Sun et al. (2019); Zhang et al. (2020) and surpassing the results achievable by each modality when used separately.
To effectively fuse the useful elements of both modalities, diverse fusion techniques and training approaches have been developed. In end-to-end networks, there are implementations utilizing residual connections Li et al. (2021) and adversarial learning techniques Ma et al. (2019, 2020). In autoencoder networks, some directly employ pretrained convolutional encoders Li and Wu (2018), others integrate fusion in different frequency domains Zhao et al. (2021), and there are also methods combining transformers with convolutional architectures Zhao et al. (2023b). These methods have preserved visual texture information to some extent. However, the resulting fused images may struggle to retain high-level semantic visual information adequately, which can lead to the obscurity of high-level visual targets and potentially suboptimal performance in downstream tasks.
@@
In an effort to learn high-level semantic information, researchers have employed downstream tasks to drive fusion networks to better learn the high-level semantic information of target objects. Some have utilized semantic segmentation tasks Tang et al. (2022b), while others have employed object detection Liu et al. (2022). However, using separate modules for image fusion and downstream tasks, and training directly with limited annotated data, makes a challenge in integrating high-level visual information into the fusion module through downstream tasks. Consequently, researchers proposed joint training of the encoder with downstream segmentation tasks Tang et al. (2023), enhancing the encoder’s ability to extract high-level visual information under limited data. There are also proposals to use meta-learning training methods Zhao et al. (2023a), allowing the model to progressively learn relevant object information and fully utilize the information brought by downstream tasks. However, these methods also introduce complexities in model architecture and training approaches, and still face the issue of insufficient labeled data, leading to a risk of overfitting on training data.
@@
Inspired by the observation that MAE encoder He et al. (2022), trained self-supervisedly on a large image dataset, has demonstrated proficiency in assimilating both low-level and high-level visual information Li et al. (2022); Liu et al. (2023a), making it an ideal choice for extracting robust features for better fusion. In this paper, we propose to leverage pretrained MAE encoder to extract the features of infrared and visible images for better fusion. By applying the pretrained model with strong representation ability, we can expect to reduce the dependence on data of IVIF task, where the paired samples are usually scarce. However, as shown in Figure 1, fusing infrared and visible light features using conventional cross-attention Li and Wu (2024); Liu et al. (2023b) may converge to the local optima, where fused image have obvious block effects and inharmonious regions. To this end, in this paper we further introduce a guided training strategy to mitigate this issue. This strategy is employed to expedite the alignment of the fusion layer output with the encoder feature domain, thereby avoiding the potential risk of converging to local optima. By adopting this approach, a simple yet effective fusion network was developed, which successfully extracts and retains both low-level image reconstruction and high-level visual features. Our contributions can be distilled into two main aspects:
@@
  • Employing a pretrained encoder, i.e., MAE, as the encoder for fusion tasks enables the fusion network to acquire comprehensive low-level and high-level visual information. This approach addresses the issue of lacking high-level visual information in fusion features and also simplifies the overall network structure. Additionally, it resolves the problem of insufficient labeled data during training.
@@
  • A guided training strategy is proposed, aiding the fusion layer in rapidly adapting to and aligning with the encoder’s feature space. This effectively resolves the issue of fusion training in vision transformer (ViT) architectures becoming trapped in local optima.
Figure 2: Workflow of our proposed MaeFuse. The upper part describes the overall architecture of the network, while the lower part elaborates on the content of the CFM and MFM structures. CFM cross-learning retains useful information, and MFM fuses enriched detail information based on the output content of CFM as a reference.
2 Related Works
@@
2.1 Deep Learning-based IVIF
The development of deep learning has brought many novel methods to image fusion. The architecture of infrared and visible image fusion networks based on deep learning can mainly be divided into autoencoder networks and end-to-end networks. Autoencoder networks can be further divided into three parts: the encoder, fusion layer, and decoder. Li et al. proposed the first pretrained fusion model Li and Wu (2018), introducing dense connections in the encoder. However, it adopted manually designed fusion strategies, which limit its flexibility and fusion performance. Later, Xu et al. used a classification saliency-based rule for image fusion Xu et al. (2021), allowing the fusion strategy to be learned as well. End-to-end fusion frameworks eliminate the need for manually designed fusion strategies. In this context, Xu et al. proposed a method Xu et al. (2020a) for general fusion tasks, which can automatically estimate the importance of source images. In addition, there are some methods that use multi-scale fusion, which also achieves very good performance Li et al. (2021). Since the image fusion problem usually does not have ground truth, some researchers have turned to exploit new training strategies, such as generative adversarial methods Ma et al. (2019, 2020); Li et al. (2020) and unsupervised methods Jung et al. (2020); Cheng et al. (2023). All the methods above mainly focus on integrating the complementary information of the two modalities and enhancing the information in the fused images. A good visual result is their common pursuit. However, they generally cannot preserve the high-level visual information of the fused images to help us highlight the interested objects.
2.2 Downstream Task-Driven IVIF
@@
To enhance the acquisition of high-level visual information for improved image fusion, Tang et al. initially employed semantic segmentation tasks as drivers for image fusion Tang et al. (2022b). In a similar vein, Liu et al. utilized object detection for the same purpose Liu et al. (2022). However, the limitation of labeled fusion datasets hampers the ability to provide semantic feedback to the fusion network through backpropagation. Addressing this, the work of Tang et al. (2023) introduces a technique to infuse high-level semantic information at the feature level within the fusion network. This is executed by co-training the encoder with both the downstream and fusion tasks. On the other hand, Zhao et al. recommended the adoption of meta-learning techniques Zhao et al. (2023a), allowing the fusion network to progressively assimilate both low-level and high-level visual information. These methods have brought fresh air to the field of image fusion, and have recently become one of the research hotspots. Nonetheless, these methods have led to increased complexity in both the fusion network and its training strategy, thereby complicating the design and implementation of fusion methods.
2.3 Pretrained Models for Feature Representation
@@
A large pre-trained feature model achieves strong generalization and performance for downstream tasks, benefiting low-sample scenarios. For pretrained models, there are mainly two types of architectures: based on convolution and based on transformers. Within the realm of traditional convolution, VGG Simonyan and Zisserman (2015) initially served as the foundational backbone for encoder architectures. Subsequently, the advent of ResNet He et al. (2016) addressed the issue of gradient vanishing, enabling deeper network structures for enhanced feature extraction. However, it’s important to note that both VGG and ResNet depend heavily on labeled data. To improve feature extraction capability and reduce dependence on labels, SimCLR Chen et al. (2020) uses contrastive learning to help the model obtain useful image representations. Regarding ViT, two primary architectures stand out: SimMIM Xie et al. (2022) and MAE He et al. (2022). Both employ a masking strategy for self-supervised training during their pretraining phase. Notably, the MAE pretrained encoder excels in feature extraction, effectively bridging low-level and high-level visual information. Consequently, in our MaeFuse framework, we leverage the MAE pretrained encoder to enrich and enhance the comprehensiveness of feature information in our fusion tasks.
3 Method
@@
In this section, we first clarify how to utilize the MAE’s pretrained encoder for feature fusion, and then provide a detailed exposition of the guided training strategy, focusing on two aspects: the loss function and the training process.
3.1 Network Architecture

### v2

@@ Abstract
In this paper, we introduce MaeFuse, a novel autoencoder model designed for Infrared and Visible Image Fusion (IVIF). The existing approaches for image fusion often rely on training combined with downstream tasks to obtain high-level visual information, which is effective in emphasizing target objects and delivering impressive results in visual quality and task-specific applications. Instead of being driven by downstream tasks, our model called MaeFuse utilizes a pretrained encoder from Masked Autoencoders (MAE), which facilities the omni features extraction for low-level reconstruction and high-level vision tasks, to obtain perception friendly features with a low cost. In order to eliminate the domain gap of different modal features and the block effect caused by the MAE encoder, we further develop a guided training strategy. This strategy is meticulously crafted to ensure that the fusion layer seamlessly adjusts to the feature space of the encoder, gradually enhancing the fusion performance. The proposed method can facilitate the comprehensive integration of feature vectors from both infrared and visible modalities, thus preserving the rich details inherent in each modal. MaeFuse not only introduces a novel perspective in the realm of fusion techniques but also stands out with impressive performance across various public datasets. The code is available at https://github.com/Henry-Lee-real/MaeFuse.
Index Terms:
Image fusion, Vision Transformer, Masked Autoencoder, Guided training
I Introduction
@@
Multimodal sensing technology has significantly contributed to the widespread use of multimodal imaging in various fields, and the fusion of infrared and visible image is the most common application technique. Infrared and Visible Image Fusion (IVIF) can merge the complementary information of these two modalities. Infrared images display thermal contours and are unaffected by lighting effects, but lacks texture and color accuracy. Visible images can capture texture but depend on lighting. Fused images can combine their strengths, and thus enhancing visual clarity and the effectiveness of downstream visual tasks [1, 2] and surpassing the results achievable by each modality when used separately.
To effectively fuse the useful elements of both modalities, diverse fusion techniques and training approaches have been developed. Traditional image fusion methods focus on enhancing visual effects by implementing various strategies, including subspace transforms [3], multi-scale transforms [4], sparse representations [5], and saliency analysis [6]. With the rise of deep learning, researchers have crafted deep learning models to enhance the visual quality of fusion outcomes. The design of architectures is primarily divided into two categories: directly concat infrared and visible images for fusion (pre-fusion), and first extract infrared and visible features separately before fusion (post-fusion). In the first approach, there are implementations utilizing residual connections [7] and adversarial learning techniques [8, 9]. In the second approach, some directly employ pretrained convolutional encoders [10], some integrate fusion in different frequency domains [11], some use Transformers as the backbone network [12, 13], and there are also some methods combining Transformers with convolutional architectures [14]. These approaches have retained some degree of visual texture information. However, the fused outcomes often display a particular style, such as increased brightness or prominent gradient features. Additionally, in certain situations, they lack the ability to make adaptive selections and adjustments, such as dimming illumination in specific areas. Fundamentally, object information, also referred to as high-level visual information, is not adequately captured and is therefore overlooked during fusion. This insufficient capture of high-level semantic cues can lead to ambiguities in identifying visual targets and potentially result in suboptimal performance in downstream tasks.
Figure 1: Diagram illustrating image-level downstream task-driven fusion and feature-level downstream task-driven fusion.
@@
In an effort to learn high-level semantic information, researchers have employed downstream tasks to drive fusion networks to better learn the high-level semantic information of target objects. Some have utilized semantic segmentation tasks [15, 16], while others have employed object detection [17, 18]. However, in the aforementioned methods, the fusion modules and downstream task modules within the network architecture do not undergo effective joint learning. This is because these methods typically feed the fused images into the downstream task models for processing. During the gradient backpropagation process, gradients first flow from the downstream task network, then through the decoder, and finally reach the fusion network. This sequence causes the gradient information to deviate to some extent. Additionally, training directly on limited annotated data makes it challenging to integrate high-level visual information into the fusion module through downstream tasks. Therefore, researchers proposed directly using the fused features as the features extracted by the downstream task network for subsequent operations [19]. This approach allows for more accurate gradient backpropagation and better performance, thereby enhancing the encoder’s ability to extract high-level visual information under limited data conditions. Fig. 1 provides a more intuitive illustration of the differences between the two aforementioned architectures. There are also suggestions to utilize meta-learning training methods [16], allowing the model to progressively learn relevant object information and fully utilize the information brought by downstream tasks. However, these methods still introduced complexities in model architecture and training approaches, and still face the issue of insufficient labeled data, leading to a risk of overfitting on training data.
Based on previous methods, the issues can generally be summarized as follows: 1) There is a deficiency of extensive paired datasets with labels for downstream tasks; 2) The need to inject high-level semantic information into the model to enhance both the visual fusion effects and performance of downstream tasks. To address the issue of scarce data, we employed pretrained models to extract visual features, as these models possess rich visual priors that can effectively capture key visual information. Furthermore, to inject high-level semantic information, we decided to use pretrained networks that already contain high-level semantic information, allowing us to naturally preserve high-level semantic information within the visual features.
@@
With the development of deep learning and pretraining models, a plethora of visual pretraining models such as MAE [20], Diffusion Model [21], and Vision Mamba [22] have emerged recently. For the Diffusion Model, although there are optimizations based on the Expectation-Maximization algorithm like DDFM [23] and prompt-based approaches like Diff-IF [24], they face practical limitations due to speed issues associated with sampling. Moreover, the images restored through sampling exhibit significant deviations from real images, such as distorted text information. As for Vision Mamba based methods [25, 26, 27], it is unclear whether this model can simultaneously capture high-level semantic and low-level texture information. MAE [20], an autoregressive pretrained model, effectively extracts both high-level [28, 29] and low-level [30] visual information, making it an ideal choice for robust feature extraction to achieve better fusion. In this paper, we propose using the pretrained MAE encoder to extract features from infrared and visible images for better fusion. Because MAE is trained via self-supervised learning, it has a strong ability to extract both high-level and low-level visual information. As a result, the fusion outcomes can maintain low-level texture and color information while also preserving robust high-level object information. We refer to these visual features that cover both high-level and low-level aspects as ’omni features’. By employing a pretrained model with strong representational capabilities, we can expect to reduce the dependency on data in the IVIF task, where paired samples with downstream task labels are typically scarce.
@@
However, as shown in Fig. 2, fusing infrared and visible features using standard cross-attention [31, 32], may result in the loss of a substantial amount of detailed information. Consequently, the fused images exhibit obvious block effects and inharmonious regions. The block effect refers to the phenomenon of discontinuity in color and texture between patches due to the ViT architecture dividing the image into patches for learning. To this end, in this paper we further introduce a guided training strategy to mitigate this issue. This strategy is employed to expedite the alignment of the fusion layer output with the encoder feature domain, thereby avoiding the potential risk of converging to local optima. By adopting this approach, a simple yet effective fusion network is developed, which successfully extracts and retains both low-level image reconstruction and high-level visual features. Our contributions can be distilled into two main aspects:
@@
  • Employing a pretrained encoder, i.e., MAE, as the encoder for fusion tasks enables the fusion network to preserve comprehensive low-level and high-level visual information. This approach addresses the issue of lacking high-level visual information in fusion features and also simplifies the overall network structure. Additionally, it also alleviate the problem of insufficient labeled data during training.
@@
  • A guided training strategy is proposed, aiding the fusion layer in rapidly adapting to and aligning with the encoder’s feature space. This effectively resolves the issue of fusion training in vision Transformer [33] (ViT) architectures becoming trapped in local optima.
The remainder of this paper is organized as follows. Section II reviews deep learning-based fusion methods, current downstream task-driven fusion approaches, and introduces some basic pretrained network frameworks. Section III provides detailed descriptions of the various design details of our work and the relevant theoretical derivations. Section IV presents both quantitative and qualitative analysis results of MaeFuse on public datasets, along with ablation studies and extended analyses of our method’s performance. Finally, Section V summarizes our approach and offers a forward-looking perspective.
II Related Work
@@
II-A Deep Learning-based IVIF
The development of deep learning has brought many novel methods to image fusion. Infrared and visible image fusion methods based on deep learning can generally be divided into post-fusion frameworks and pre-fusion frameworks.
For the post-fusion framework, models are often pre-trained using large-scale datasets, so that the encoder obtains good image feature information. Li et al. firstly introduced the pretrained fusion model known as DenseFuse [10], which is comprised of three parts: an encoder layer, a fusion layer, and a decoder layer. In the fusion layer, they applied strategies such as element-wise addition or $\ell_{1}$-norm, and implemented dense connections within the encoder layer to achieve effective fusion outcomes. Later, they also developed a multi-scale architecture and nested connections to enhance the extraction of more detailed features [34] [7]. Zhao et al. created a novel encoder designed for multi-scale decomposition [11], aimed at extracting both detailed and background features. In a similar vein, Tang et al. applied Retinex theory to improve nighttime image fusion results by employing several encoders [35], which separate the illumination and reflection components of visible images. However, the methods mentioned above adopted manually designed fusion strategies, which limit their flexibility and fusion performance. Subsequently, Xu et al. used a classification saliency-based rule for image fusion [36], allowing the fusion strategy to be learned as well. Recently, some teams have explored fusion using more information scales based on this structure [37], avoiding the shortcomings of only dividing information into low-frequency and high-frequency. Additionally, other teams have incorporated textual information control at the feature level [38], making the fusion process more controllable.
Pre-fusion frameworks eliminate the need for manually designed loss functions, network architectures, and learning paradigms. Since fused images lack ground truth, researchers have developed various loss functions based on the characteristics of image fusion to represent the most basic fusion information, thereby providing targeted guidance for training. Based on common intensity and gradient losses, Ma et al. designed a fusion loss based on a significant target mask, used to selectively fuse target and background areas [39].
Moreover, considering variations in illumination, they designed an illumination-aware loss function [40]. Structural similarity loss [41] and perceptual loss [42] were also introduced to restrict fusion outcomes and avoid distortion of structural information while ensuring good visual perception. Beyond loss functions, numerous innovative network architectures have been developed, such as residual blocks [39], aggregated residual dense blocks [41], and gradient residual dense blocks [15], along with fusion modules including cross-modality differential aware fusion modules [40], interaction fusion modules [43], global spatial attention modules [44], and biphasic recurrent fusion modules [45], to ensure the comprehensiveness of information in the fusion results. Researchers have also introduced some new learning paradigms, such as generative adversarial mechanisms and unsupervised learning. Ma et al. were the first to introduce generative adversarial networks into the field of image fusion with the development of FusionGAN [8]. This model uses a discriminator to force the generator to retain more texture details from visible images, although it does not adequately consider information from infrared images. The successors to FusionGAN, such as DDcGAN [9], AttentionFGAN [46], and SDDGAN [47], have designed dual discriminators to address this issue, avoiding the modal imbalance caused by a single discriminator. Unsupervised learning methods [48, 49] are also employed. Some teams have proposed FusionBooster [50] to compensate for the loss of detailed content in fused images, enabling the fusion results to recover relevant texture information.
In recent years, due to the Transformer’s superior long-range learning capability, the Vision Transformer (ViT) has gradually replaced CNN as the basic framework for vision. As a result, some Transformer-based fusion models, such as IFT [51], AFT [52], SwinFuse [13],and SwinFusion [12], have been developed to fully explore the long-range dependencies in source images. Considering that infrared and visible light images often exhibit various degrees of misalignment in practical applications, the latest methods (such as RFNet [53], UMF-CMGR [43], ReCoNet [45], and SuperFusion [44]) incorporate an alignment module before the fusion module to correct misalignments in the source images. Additionally, some methods, including PMGI [54], IFCNN [55], U2Fusion [56]and DeFusion [57], handle various image fusion tasks in a unified manner, as there are commonalities among these tasks. Particularly, U2Fusion [56] trains a unified model to handle multiple fusion tasks, promoting cross-fertilization among different fusion tasks. All the methods above mainly focus on integrating the complementary information of the two modalities and enhancing the information in the fused images. A good visual result is their common pursuit. However, they generally cannot preserve the high-level visual information of the fused images to help us highlight the interested objects.
II-B Downstream Task-Driven IVIF
@@
To enhance the acquisition of high-level visual information for improved image fusion, Tang et al. initially employed semantic segmentation tasks as drivers for image fusion [15]. In a similar vein, Liu et al. utilized object detection for the same purpose [17]. However, the limitation of labeled fusion datasets hampers the ability to provide semantic feedback to the fusion network through backpropagation. Addressing this, the work of [19] introduces a technique to infuse high-level semantic information at the feature level within the fusion network. This is executed by co-training the encoder with both the downstream and fusion tasks. On the other hand, Zhao et al. recommended the adoption of meta-learning techniques [16], allowing the fusion network to progressively assimilate both low-level and high-level visual information. Meanwhile, to better utilize image segmentation information, Liu et al. injected features from the segmentation task back into the fusion network [58], guiding the image to more effectively integrate segmentation information. Recently, Zhang et al. combined object detection with diffusion models [59], aiming to achieve better fused images by leveraging the object detection’s ability to capture information and the powerful image generation capabilities of diffusion models. These methods have brought fresh air to the field of image fusion, and have recently become one of the research hotspots. Nonetheless, these methods have led to increased complexity in both the fusion network and its training strategy, thereby complicating the design and implementation of fusion methods.
II-C Pretrained Models for Feature Representation
@@
A large pre-trained feature model achieves strong generalization and performance for downstream tasks, benefiting low-sample scenarios. For pretrained models, there are mainly two types of architectures: based on convolution and based on Transformers. Within the realm of traditional convolution, VGG [60] initially served as the foundational backbone for encoder architectures. Subsequently, the advent of ResNet [61] addressed the issue of gradient vanishing, enabling deeper network structures for enhanced feature extraction. However, it’s important to note that both VGG and ResNet depend heavily on labeled data. To improve feature extraction capability and reduce dependence on labels, SimCLR [62] uses contrastive learning to help the model obtain useful image representations. Regarding ViT, two primary architectures stand out: SimMIM [63] and MAE [20]. Both employ a masking strategy for self-supervised training during their pretraining phase. Notably, the MAE pretrained encoder excels in feature extraction, effectively bridging low-level and high-level visual information. Consequently, in our MaeFuse framework, we leverage the MAE pretrained encoder to enrich and enhance the comprehensiveness of feature information in our fusion results.
III Method
@@
In this section, we first use mathematical derivations to demonstrate that the encoder requires high-level semantic information. Then, we clarify how to utilize the MAE’s pretrained encoder for feature fusion. Finally, we provide a detailed exposition of the guided training strategy, focusing on two aspects: the loss function and the training process.
Figure 3: Workflow of our proposed MaeFuse. The upper part describes the overall architecture of the network, while the lower part elaborates on the content of the CFM and MFM structures. CFM cross-learning retains useful information, and MFM fuses enriched detail information based on the output content of CFM as a reference. The FFN employs merely two layers of fully connected neural networks, aimed specifically at enhancing the model’s capacity for non-linear learning.
III-A Proposition
First, we aim to use formulas to demonstrate that only when the encoder obtains sufficient semantic information will our final fusion result highlight adequate semantic information. We define the contents of the labels: $V_{f}$ represents the features corresponding to the visible image obtained by the encoder, $I_{f}$ represents the features corresponding to the infrared image obtained by the encoder, and $\hat{F}$ represents the fusion result obtained by the existing network. The result fused through the existing network is a definite value, with the probability of 1,
$P(\hat{F}|V_{f},I_{f})=1.$ (1)
We aim to establish the correlation between the features obtained from the two modalities through the encoder and the global high-level semantic information. Therefore, we define $S$ to represent the inherent high-level semantic information of the two modalities. The following formula represents the probability of obtaining the inherent high-level semantic information of the two modalities given the infrared and visible image features,
$P(S|V_{f},I_{f}).$ (2)
[…]
$P(S|V_{f},I_{f})$ $=P(S,\hat{F}|V_{f},I_{f})+P(S,\neg\hat{F}|V_{f},I_{f}),$
$P(S,\neg\hat{F}|V_{f},I_{f})$ $=P(S|\neg\hat{F},V_{f},I_{f})(1-P(\hat{F}|V_{f},I_{f})),$
$P(S|V_{f},I_{f})$ $=P(S,\hat{F}|V_{f},I_{f}),$ (3)
$P(S,\hat{F}|V_{f},I_{f})$ $=P(S|\hat{F},V_{f},I_{f})P(\hat{F}|V_{f},I_{f}),$
$P(S,\hat{F}|V_{f},I_{f})$ $=P(S|\hat{F},V_{f},I_{f}),$
$P(S,\hat{F}|V_{f},I_{f})$ $=P(S|\hat{F}).$ (4)
From the above equations, we can obtain the following:
$P(S|V_{f},I_{f})$ $=P(S,\hat{F}|V_{f},I_{f})=P(S|\hat{F})\leq P(S|F).$ (5)
Where $F$ represents the optimal fusion effect, and thus, the optimal fusion effect undoubtedly preserves the most high-level semantic information.
By combining the above equations, we can consider that to achieve the best fusion result, the encoder must effectively capture high-level semantic information from the image features. Furthermore, this information must be efficiently utilized within the fusion structure to attain a satisfactory fusion outcome.
III-B Network Architecture